The random variable X, representing the number of errors per 100 lines of software code, has the following probability distribution:

x                              2                             3                              4                             9                            

f(x)                         0.1                          0.2     c                            0.1                         

where c is defined as the probability of having 4 errors.

What would be the expected value of X ?
Answer #1

Sum of probabilities equal to 1

f(x) = 1

0.1 + 0.2 + C + 0.1 = 1

C + 0.4 = 1

C = 0.6

E(X) = X * f(X)

= 2 * 0.1 + 3 * 0.2 + 4 * 0.6 + 9 * 0.1

= 4.1